# LERa: Replanning with Visual Feedback in Instruction Following

### v1

@@ Abstract
Large Language Models are increasingly used in robotics for task planning, but their reliance on textual inputs limits their adaptability to real-world changes and failures. To address these challenges, we propose LERa – Look, Explain, Replan – a Visual Language Model-based replanning approach that utilizes visual feedback. Unlike existing methods, LERa requires only a raw RGB image, a natural language instruction, an initial task plan, and failure detection – without additional information such as object detection or predefined conditions that may be unavailable in a given scenario. The replanning process consists of three steps: (i) Look, where LERa generates a scene description and identifies errors; (ii) Explain, where it provides corrective guidance; and (iii) Replan, where it modifies the plan accordingly. LERa is adaptable to various agent architectures and can handle errors from both dynamic scene changes and task execution failures. We evaluate LERa on the newly introduced ALFRED-ChaOS and VirtualHome-ChaOS datasets, achieving a 40% improvement over baselines in dynamic environments. In tabletop manipulation tasks with a predefined probability of task failure within the PyBullet simulator, LERa improves success rates by up to 67%. Further experiments, including real-world trials with a tabletop manipulator robot, confirm LERa’s effectiveness in replanning. We demonstrate that LERa is a robust and adaptable solution for error-aware task execution in robotics. The code is available at https://lera-robo.github.io.
@@ I INTRODUCTION
However, there is no guarantee that a robot will successfully execute a generated task plan in a real environment. This is because LLMs rely on common sense knowledge acquired during training and do not account for dynamic environmental changes or task failures. For example, given the instruction “Heat a slice of pizza,” the model might generate a plan that includes the task “Open the microwave.” However, if the microwave is already open – perhaps because the previous user forgot to close it – the plan may result in an error, preventing successful execution. Similarly, if the slice of pizza falls while being placed inside, the robot may continue executing the plan without recognizing the failure, leading to an incomplete or incorrect outcome.
@@
To address this problem, various LLM-based replanning approaches have been proposed [8, 9, 10]. However, LLMs rely solely on textual input, which does not fully capture the state of the environment and requires this state to be represented in textual form, such as a list of objects or a scene description [2, 11, 5, 6, 8]. This textual representation is typically derived from the output of computer vision systems, such as object detectors or semantic segmenters. Visual Language Models (VLMs), on the other hand, can directly process both visual and textual modalities, making them more suitable for capturing the state of the environment and enabling replanning. However, most modern VLM-based replanning approaches: (i) are designed for a specific task planner or agent architecture, making them difficult to adapt to different architectures [12, 13, 14, 15], (ii) require not only information from the RGB camera and the instruction with the corresponding task plan for replanning but also additional data, such as pre- and post-action conditions, target object positions, and object bounding boxes – information that may not always be available in a given scenario [13, 14], (iii) are primarily designed to handle a single type of error, limiting their applicability to other error types [12, 13, 15].
@@
Figure 1: An approach without error handling can be enhanced with the LERa module to dynamically update the queue of subtasks when errors occur during plan execution. On the left (solid arrows), the plan is generated once at the beginning of an episode and remains unchanged throughout. With LERa (dashed arrows), when an error occurs during the execution of the current subtask, a Visual Language Model is used to replan based on visual feedback.
@@
In this paper, we address these problems by proposing LERaLook, Explain, Replan – a VLM-based approach for replanning with visual feedback. LERa uses only basic information for replanning: (i) a raw RGB camera image, without requiring detected objects, masks, or any other additional information, (ii) information about action execution failure, without needing an explanation of the reason for the failure, (iii) the initial natural language instruction, and (iv) the task plan. LERa makes no specific assumptions about the agent’s architecture and can extend any agent that detects action execution failure and requests replanning (see Fig. 1). The replanning process in LERa consists of three steps. Look – given the raw image from the RGB camera and information about the action failure, LERa generates a scene description and identifies the cause of the error. Explain – based on the identified information, LERa explains how to correct the plan and addresses the replanning problem without being constrained to a specific output format. Replan – LERa modifies the old plan by incorporating the proposed changes and passes the final version of the plan to the agent for execution. LERa is capable of handling errors from dynamic environmental changes and action execution failures.
@@
To evaluate LERa’s ability to replan we collected the ALFRED-ChaOS and VirtualHome-ChaOS datasets, each containing more than 500 episodes, based on ALFRED [16] and VirtualHome [17] environments. LERa improves success rates by over 40% compared to baselines, including a one-step VLM-based replanner. We also test LERa on 10 tabletop manipulation tasks in PyBullet [18] with stochastic action failures. LERa achieves up to 67% higher success than a no-replanning baseline.
@@
We also conducted extensive ablation experiments, confirming the necessity of all three steps in the replanning process. Additionally, we performed experiments with different VLMs, demonstrating how VLM quality affects replanning performance. Further experiments with a non-ideal subtask checker on ALFRED-ChaOS show the consistent superiority of LERa when working with imperfect error detection. To demonstrate the generalizability of our approach to real-world scenarios, we integrated LERa into the control system of a real robot for tabletop manipulation. LERa successfully handled replanning in 15 out of 18 trials.
@@
    We introduce three different environments, based on well-known simulators, that can be used to evaluate replanning approaches.
@@ II RELATED WORK
VLM-based approaches are widely used in robotics for replanning, but they usually either rely on additional information beyond the camera image, assume a specialized agent architecture, or replan only for a specific type of error. In [19], the VLM is used as the initial task planner, and the replanning problem is not structurally separated. Although this approach significantly simplifies the architecture, it does not account for the specific features of the environment or the robot in the initial planning [13]. In [15], a complex interaction structure between planning modules at different levels is used, where VLM is first employed to confirm the correctness of an action choice before its execution in the environment. In [12], replanning is framed as a visual question-answering problem, handling only errors related to action failures and requiring information about the preconditions and effects of actions. In [14], a VLM-based Extra Bot is used as an action failure detector by analyzing frames before and after an action is performed. Replanning is handled by the Design Bot, which also functions as a task planner, imposing constraints on the agent’s architecture. In [13], VLM is used to generate intermediate goals instead of actions during replanning. In this case, information about detected objects, their relationships, and collisions is provided to VLM along with the image. In [20], VLM is further trained to detect and explain errors.
@@ III-A Problem Formulation
We consider the problem of replanning as the problem of adjusting the initial plan $P$, which consists of a sequence of actions (tasks) $P=(a_{1},a_{2},\dots,a_{n})$ (where $n$ is a plan length) executable in the environment, upon receiving information about the occurrence of an error during the execution of an action (task). We do not consider low-level actions – those resulting from motion planning rather than task planning – as components of the plan. Therefore, we use the terms “action” and “task’’ synonymously where this does not cause confusion. By plan adjustment, we mean modifying the sequence of actions in the plan $P=(a_{1},a_{2},\dots,a_{n})$ to $P^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{n^{\prime}})$ so that execution can successfully continue in the environment. Since we do not address the problem of initial planning, we assume that the plan $P$ always leads to achieving the goal in the absence of environmental changes or unsuccessful action execution.
We assume that the replanner $R$ operates only on information available after the planning stage – namely, natural language instructions $I$, the task plan $P$, and information about the occurrence of an error at time $t$ (denoted as $E_{t}$), obtained using the Subtask Checker $SC$. Additionally, $R$ has access to the current observation $O_{t}$ from an RGB camera but does not use any other information. Thus, the problem is reduced to creating a function $R$ that, given all the described inputs, generates a corrected plan $P^{\prime}$: $R(O_{t},E_{t},I,P)=P^{\prime}$.
@@ III-B The LERa Approach
LERa makes fairly general assumptions about an agent’s architecture, requiring three main modules: (1) Task Planner, (2) Subtask Executor, and (3)Subtask Checker (Fig.3). The Task Planner (1) generates a task plan $P$ from language instructions $I$ and additional information, managing task execution by assigning the next action $a_{i}$. The Subtask Executor (2) produces low-level actions, potentially incorporating navigation and motion planning. The The Subtask Checker (3) verifies task execution using environment feedback, requesting the next action $a_{i+1}$ if successful or triggering replanning from LERa (4) if an error occurs. Due to its modular structure, LERa does not depend on how the initial plan is generated. It requires only that the input to the replanner be a sequence of actions to revise, enabling integration with various high-level planning systems.
The LERa module updates the agent’s plan $P$ to a revised plan $P^{\prime}$, mitigating execution errors. It either sends $P^{\prime}$ to the Task Planner or directly provides a corrected action $a^{\prime}_{i+1}$ to the Subtask Executor. LERa operates in three main steps: Look, Explain, and Replan. The algorithm for LERa’s operation is presented in Algorithm 1, and Fig. 2 illustrates a step-by-step example of replanning.
@@
Look Step: At the moment of task $a_{t}$ execution failure ($SC(a_{t})=0$), the agent captures the visual observation $O_{t}$. Based on $O_{t}$ and $a_{t}$, LERa creates a scene description and provides examples of various failure reasons. The predicted description is called $L$.
@@
The main experiments were performed with Oracle versions of the agent modules, which are capable of executing high-level actions with minimal errors, as our goal was to isolate LERa’s replanning performance from errors in other modules.
@@
4:Current task plan $P=(a_{t},...,a_{n})$,
5:Failed action $a_{t}$
6:Updated action plan $P^{\prime}$ that ensures successful task completion
7:$L\leftarrow\text{Look}(o_{t},a_{t})$triangleright Generate scene description and identify failure cause
8:$E\leftarrow\text{Explain}(I,L,P)$triangleright Explain error and propose conceptual plan
9:$P^{\prime}\leftarrow\text{Replan}(I,E,P)$triangleright Predict updated sequence of actions
10:return $P^{\prime}$
Figure 3: The proposed structure of an agent to use LERa consists of three modules: (1) Task Planner, (2) Subtask Executor, and (3) Subtask Checker. Blue is for language, yellow is for planning and execution, red is for self-checking, and green is for replanning.
@@ III-C Prompting Strategies
The LERa method utilizes three different prompt templates for each step to solve specific tasks (Fig. 4). In the Look Step, LERa defines the reasons for the error and provides simple ideas for a solution. To prevent the VLM from being overloaded with unnecessary information, only a limited set of key facts about the virtual environment is included in the system instruction. For the same reason, only the first step of the current plan is provided to the model along with the observation. The Explain Step is designed to address the replanning task without being constrained by any output format. The objective of this stage is to reason about the given task instruction, the existing plan, and the newly acquired visual information in order to generate an updated task plan. In addition, the model is required to provide an explanation for each predicted step, ensuring that the output contains as much relevant information as possible to support the final replanning step. The Replan Step constructs the final corrected task plan. Its prompt includes few-shot examples to constrain formatting, a list of available environment actions with descriptions, the existing task plan, and the newly predicted task plan with step explanations. This step ensures the new plan is executable in the virtual environment, as the Explain Step may produce infeasible actions. Additionally, the Replan Step refines the plan to align with the environment or correct logical errors.
@@ IV Experiments
Figure 5: We evaluate LERa across diverse set of environments. From left to right, visual observations from: ALFRED, VirtualHome, TableTop-PyBullet, and TableTop-RoboticStand.
To demonstrate the generalizability of the LERa approach, we evaluated it in diverse simulation environments and on a real table-top manipulation robot (Fig. 5). We conducted experiments with different parts of LERa to assess the contribution of each component to overall success. Additional evaluations focused on measuring the impact of the VLM on the overall success rate. Moreover, we tested LERa with a non-ideal checker to evaluate its ability to correctly assess the environment state and restore the original plan when no actual errors occurred. In addition, we conducted experiments on a real robot to demonstrate the applicability of the method in real world conditions.
@@ IV-A Environments
We conducted diverse set of experiments in three virtual environments: ALFRED [16], VirtualHome [17] and PyBullet. In ALFRED and VirtualHome, experiments focused on assessing the approach’s ability to handle errors caused by unexpected object state changes. In PyBullet, we evaluated its effectiveness in scenarios where the last executed action completed, but the intended goal was not achieved. Such tasks require the agent to detect partial failure, infer its cause, and adjust the plan accordingly
@@
ALFRED-ChaOS. In ALFRED, an embodied agent is asked to complete a household task, e.g. “Put a clean sponge on a metal rack”, by giving low-level instructions, e.g. “Go to the left and face the faucet side of the bathtub. Pick up left most …”. We took episodes from ALFRED’s “valid_seen” and “valid_unseen” splits, filtered out the repetitive ones as they corresponded to the same ground truth (GT) plans, and changed states of objects that affect the execution of the task. As a result, we obtained 251 episodes with 290 cases of objects with changed states. We combined these episodes with their original versions and ended up with a total of 502 episodes. We refer to the resulting dataset as ALFRED-ChaOS, which stands for Changed Object States.
@@
VirtualHome-ChaOS. VirtualHome is similar to ALFRED as an embodied agent is asked to complete household tasks. In this environment, we adapted three types of tasks: heating objects in the microwave, placing objects in the refrigerator, and washing objects in the dishwasher. For each of the 50 scenes, we collected two distinct tasks for each task type, involving different objects. Scenes that lacked essential containers for the tasks were excluded from the dataset. Using this approach, we created 262 tasks requiring interactions with various objects in the scene. To implement changing in objects states, we modified the states of key containers relevant to each task. Combining these modified tasks with the original ones resulted in a total of 524 tasks. We refer to the resulting dataset as VirtualHome-ChaOS.
@@
TableTop-PyBullet. In Alfred and Virtual Home we evaluated replanning capabilities in scenarios where the object’s state differed from the initial state. To evaluate the ability to resolve errors that occurred during action execution, we formed a series of experiments in PyBullet. Agent has two primary actions: ‘pick’ and ‘place’ but may drop the object during their execution with predefined probability. We have prepared 10 tasks and run each 10 times with different object’s locations and drop moments resulting in 100 scenarios. Each task has a different plan length and difficulty. For example “Place red block in red bowl” and “Build two towers in bowls: blue on red, yellow on green” have lengths of one and four pick-place action pairs. In the first task agent must perform action sequence of: “locate red block”, “pick red block”, “locate red bowl”, “place red bowl”. The second task plan is four times larger, as agent has to move four blocks instead of one.
@@
TableTop-RoboticStand. To evaluate LERa in a real-world environment, we use the XArm6 robotic platform, which is equipped with XArm grippers and RealSense L515 and D435 sensors. It performs manipulation tasks with two primary actions: pick and place. The planner, based on an open-source LLM, generates the plan using object recognition data, while the fine-tuned VLM-based checker evaluates the success of the action execution. Re-planning is triggered when checker detects execution errors, such as misplacement or misclassification of objects. The experiments consisted of 18 different tasks involving the manipulation of various sets of objects from a table to a container. Notably, in this real world setting, errors frequently occurred due to inaccuracies in the gripper position predicted by the agent. This lead to failures in grasping or object placement. Additionally, some episodes involved initial misclassification by the visual detector, resulting in an incorrect initial plan and, consequently, the inability to correctly execute the human instruction. Due to frequent errors, replanning was triggered, allowing LERa to adjust the original plan not only to correct the last observed mistake but also to ensure the successful execution of the intended human instruction.

### v2

@@ Abstract
Large Language Models are increasingly used in robotics for task planning, but their reliance on textual inputs limits their adaptability to real-world changes and failures. To address these challenges, we propose LERa — Look, Explain, Replan — a Visual Language Model-based replanning approach that utilizes visual feedback. Unlike existing methods, LERa requires only a raw RGB image, a natural language instruction, an initial task plan, and failure detection — without additional information such as object detection or predefined conditions that may be unavailable in a given scenario. The replanning process consists of three steps: (i) Look — where LERa generates a scene description and identifies errors; (ii) Explain — where it provides corrective guidance; and (iii) Replan — where it modifies the plan accordingly. LERa is adaptable to various agent architectures and can handle errors from both dynamic scene changes and task execution failures. We evaluate LERa on the newly introduced ALFRED-ChaOS and VirtualHome-ChaOS datasets, achieving a 40% improvement over baselines in dynamic environments. In tabletop manipulation tasks with a predefined probability of task failure within the PyBullet simulator, LERa improves success rates by up to 67%. Further experiments, including real-world trials with a tabletop manipulator robot, confirm LERa’s effectiveness in replanning. We demonstrate that LERa is a robust and adaptable solution for error-aware task execution in robotics. The project page is available at https://lera-robo.github.io.
@@ I INTRODUCTION
However, there is no guarantee that a robot will successfully execute a generated task plan in a real environment. This is because LLMs rely on common sense knowledge acquired during training and do not account for dynamic environmental changes or task failures. For example, given the instruction “Heat a slice of pizza,” the model may well generate a plan that includes the task “Open the microwave.” However, if the microwave is already open, perhaps because the previous user forgot to close it, the plan will result in an error, preventing successful execution. Similarly, if the pizza slice falls while being placed inside, the robot may continue executing the plan without recognizing the failure, leading to an incomplete or incorrect outcome.
@@
To address this problem, various LLM-based replanning approaches have been proposed [8, 9, 10]. However, LLMs rely solely on textual input, which does not fully capture the state of the environment and requires this state to be represented in textual form, such as a list of objects or a scene description [2, 11, 5, 6, 8]. This textual representation is typically derived from the output of computer vision systems, such as object detectors or semantic segmenters. Visual Language Models (VLMs), on the other hand, can directly process both visual and textual modalities, making them more suitable for capturing the state of the environment and enabling replanning. However, most modern VLM-based replanning approaches: (i) are designed for a specific task planner or agent architecture, making them difficult to adapt to different architectures [12, 13, 14, 15]; (ii) require not only information from the RGB camera and the instruction with the corresponding task plan for replanning but also additional data, such as pre- and post-action conditions, target object positions, and object bounding boxes — information that may not be available in a given scenario [13, 14]; (iii) are primarily designed to handle a single type of error, limiting their applicability to other error types [12, 13, 15].
@@
Figure 1: An approach without error handling can be enhanced by the LERa module to dynamically update the plan (a queue of tasks) when errors occur during plan execution. On the left (solid arrows), the plan is generated once at the beginning of an episode and remains unchanged throughout. With LERa (dashed arrows), a Visual Language Model is used to replan based on visual feedback whenever an error occurs.
@@
In this paper, we address these problems by proposing LERa — Look, Explain, Replan — a VLM-based approach for replanning with visual feedback. LERa uses only basic information for replanning: (i) a raw RGB camera image, without requiring detected objects, masks, or any other additional information; (ii) failure-triggering action, without requiring an explanation of its cause; (iii) the initial natural language instruction; and (iv) the task plan. LERa makes no specific assumptions about the agent’s architecture and can extend any agent that detects action execution failures and requests replanning (see Fig. 1). Moreover, LERa is capable of handling errors caused by dynamic environmental changes and action execution failures. The replanning process in LERa consists of three steps. Look — given the raw RGB camera image and information about the action failure, LERa generates a scene description and identifies the cause of the error. Explain — based on the identified information, LERa explains how to correct the plan and addresses the replanning problem without being constrained to a specific output format. Replan — LERa modifies the old plan by incorporating the proposed changes and passes the final version of the plan to the agent for execution.
@@
To evaluate LERa’s ability to replan, we collected the ALFRED-ChaOS and VirtualHome-ChaOS datasets, each containing more than 500 episodes, based on ALFRED [16] and VirtualHome [17] environments. LERa improves success rates by over 40% compared to baselines, including a one-step VLM-based replanner. We also test LERa on 10 tabletop manipulation tasks in PyBullet [18] with stochastic action failures. LERa achieves up to 67% higher success than a no-replanning baseline.
@@
We also conducted extensive ablation experiments that confirmed the necessity of all three steps in the replanning process. Additionally, we performed experiments with different VLMs that demonstrated how VLM quality affects replanning performance. Further experiments with a non-ideal subtask checker on ALFRED-ChaOS showed the consistent superiority of LERa when working with imperfect error detection. To demonstrate the generalizability of our approach to real-world scenarios, we integrated LERa into the control system of a real robot for tabletop manipulation. LERa successfully handled replanning in 15 out of 18 trials.
@@
    We introduce three different environments based on well-known open-source simulators, which can be used to evaluate replanning approaches.
@@ II RELATED WORK
VLM-based approaches are widely used in robotics for replanning, but they usually either rely on additional information beyond the camera image, assume a specialized agent architecture, or replan only for a specific type of error. In [19], the VLM is used as the initial task planner, and the replanning problem is not structurally separated. Although this approach significantly simplifies the architecture, it does not account for specific features of the environment or the robot in the initial planning [13]. In [15], a complex interaction structure is used between planning modules at different levels, where VLM is employed to confirm the correctness of an action choice before its execution in the environment. In [12], replanning is framed as a visual question-answering problem, handling only errors related to action failures and requiring information about the preconditions and effects of actions. In [14], a VLM-based Extra Bot is used as an action failure detector by analyzing frames before and after an action is performed. Replanning is handled by the Design Bot, which also functions as a task planner, imposing constraints on the agent’s architecture. In [13], VLM is used to generate intermediate goals instead of actions during replanning. In this case, information about detected objects, their relationships, and collisions is provided to VLM along with the image. In [20], VLM is further trained to detect and explain errors.
@@ III-A Problem Formulation
We consider the problem of replanning as the problem of adjusting an initial plan $P$, which consists of a sequence of actions (tasks) $P=(a_{1},a_{2},\dots,a_{n})$ (where $n$ is the plan length) executable in the environment, upon receiving information about the occurrence of an error during execution of an action (task). We do not consider low-level actions, which result from motion planning rather than task planning, to be components of the plan. Therefore, we use the terms “action” and “task’’ synonymously where this does not cause confusion. By plan adjustment, we mean modifying the sequence of actions in the plan $P=(a_{1},a_{2},\dots,a_{n})$ to $P^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{n^{\prime}})$ so that execution can continue successfully in the environment. Since we do not address the problem of initial planning, we assume that the plan $P$ always leads to the achievement of the goal in the absence of environmental changes or unsuccessful action execution.
We also assume that a replanner $R$: 1) operates only on information available after the planning stage — namely, a natural language instruction $I$, the task plan $P$, and information about the occurrence of an error at time $t$ (denoted as $E_{t}$ and obtained using a task checker $TC$); 2) has access to the current raw RGB observation $O_{t}$. Thus, the problem is reduced to creating a function $R$ that, given all the described inputs, generates the corrected plan $P^{\prime}$: $R(O_{t},E_{t},I,P)=P^{\prime}$.
@@ III-B The LERa Approach
LERa makes fairly general assumptions about an agent’s architecture, requiring three main modules: 1) a Task Planner, 2) a Task Executor, and 3) a Task Checker (Fig.3). After the Task Planner (1) has generated a task plan $P$ from the language instruction $I$ and, possibly, the environment initial state, the workflow is as follows: (i) the Task Planner (1) manages task execution by assigning the next action $a_{i}$; (ii) the Task Executor (2) produces low-level actions, potentially incorporating navigation and motion planning; (iii) the Task Checker (3) verifies task execution using environment feedback, then either requests the next action $a_{i+1}$ if execution is successful or triggers replanning by LERa (4) if an error occurs. Due to its modular structure, LERa does not depend on how the initial plan is generated. It requires only that the input to the replanner be a sequence of actions to revise, enabling integration with various high-level planning systems.
The LERa module updates the agent’s plan $P$ to a revised plan $P^{\prime}$, mitigating execution errors. It operates in three main steps: Look, Explain, and Replan. The algorithm for LERa’s operation is presented in Algorithm 1, and Fig. 2 illustrates a step-by-step example of the replanning process.
@@
Look Step: At the moment of the execution failure of a task $a_{t}$ ($TC(a_{t})=0$), the agent captures the visual observation $O_{t}$. Based on $O_{t}$ and $a_{t}$, LERa creates a scene description and provides various failure reasons. The predicted description is denoted by $L$.
@@
4:Current task plan $P=(a_{t},...,a_{n}).$
5:Updated action plan $P^{\prime}$ that ensures successful task completion
[…]
9:return $P^{\prime}$
Figure 3: The supposed structure of an agent to use LERa consists of three modules: (1) Task Planner, (2) Task Executor, and (3) Task Checker. Blue is for language, yellow is for planning and execution, red is for self-checking, and green is for replanning.
@@ III-C Prompting Strategies
The LERa method utilizes three different prompt templates — one for each step — to solve specific tasks (Fig. 4). In the Look Step, LERa identifies the causes of the error and provides simple ideas for a solution. To prevent the VLM from being overloaded with unnecessary information, only a limited set of key facts about the environment is included in the system instruction. For the same reason, only the first step of the current plan is provided to the model along with the observation. The Explain Step is designed to address the replanning task without being constrained by any output format. The objective of this stage is to reason about the given task instruction, the existing plan, and the newly acquired visual information in order to generate an updated task plan. In addition, the model is required to provide an explanation for each predicted step, ensuring that the output contains as much relevant information as possible to support the final replanning step. The Replan Step constructs the final corrected task plan. Its prompt includes few-shot examples to constrain formatting, a list of available environment actions with descriptions, the existing task plan, and the newly predicted task plan with step explanations. This step ensures that the new plan is executable in the virtual environment, as the Explain Step may produce infeasible actions. Additionally, the Replan Step refines the plan to align it with the environment or to correct logical errors.
@@ IV EXPERIMENTS
Figure 5: We evaluate LERa across diverse set of environments. From left to right, visual observations from: ALFRED, VirtualHome, TableTop PyBullet, and TableTop Robotic Stand.
To demonstrate the generalizability of the LERa approach, we evaluated it in diverse simulation environments and on a real tabletop manipulation robot (Fig. 5). We conducted experiments with different parts of LERa to assess the contribution of each component to overall success. Additional evaluations focused on measuring the impact of the VLM on the overall success rate. Moreover, we tested LERa with a non-ideal checker to evaluate its ability to correctly assess the environment state and restore the original plan when no actual errors occurred. In addition, we conducted experiments on a real robot to demonstrate the applicability of the method in real world conditions.
@@ IV-A Environments
We conducted a diverse set of experiments in three virtual environments: ALFRED [16], VirtualHome [17] and PyBullet. In ALFRED and VirtualHome, the experiments focused on assessing the approach’s ability to handle errors caused by unexpected object state changes. In PyBullet, we evaluated its effectiveness in scenarios where the last executed action completed but the intended goal was not achieved. Such tasks require the agent to detect partial failure, infer its cause, and adjust the plan accordingly.
@@
ALFRED-ChaOS. In ALFRED, an embodied agent is asked to complete a household task, e.g. “Put a clean sponge on a metal rack”, by following low-level instructions, e.g. “Go to the left and face the faucet side of the bathtub. Pick up left most …”. We took episodes from ALFRED’s “valid_seen” and “valid_unseen” splits, filtered out the repetitive ones as they corresponded to the same ground truth (GT) plans, and changed states of objects that affect the execution of the task. As a result, we obtained 251 episodes with 290 cases of objects with changed states. We combined these episodes with their original versions and ended up with a total of 502 episodes. We refer to the resulting dataset as ALFRED-ChaOS, which stands for Changed Object States.
@@
VirtualHome-ChaOS. VirtualHome is similar to ALFRED as an embodied agent is asked to complete household tasks. In this environment, we adapted three types of tasks: heating objects in a microwave, placing objects in a refrigerator, and washing objects in a dishwasher. Two distinct tasks corresponding to each task type were collected for each of the 50 scenes, ensuring variation in the involved objects. Then, scenes that lacked essential containers for the tasks were excluded from the set. Using this approach, we created 262 tasks requiring interactions with various objects in the scene. To implement changing in objects states, we modified the states of key containers relevant to each task. Combining these modified tasks with the original ones resulted in a total of 524 tasks. We refer to the resulting dataset as VirtualHome-ChaOS.
@@
TableTop PyBullet. In ALFRED and VirtualHome, we evaluated replanning capabilities in scenarios where the objects’ states differed from the initial states. To evaluate the ability to resolve errors that occur during action execution, we designed a series of experiments in PyBullet. The agent has two primary actions: “pick” and “place”, but may drop the object during their execution with a predefined probability. We prepared 10 tasks and run each 10 times with different object’s locations and drop moments, resulting in 100 scenarios. Each task has a different plan length and difficulty. For example, “Place red block in red bowl” and “Build two towers in bowls: blue on red, yellow on green” require 4 and 16 actions, respectively. In the first task, the agent must perform the actions “locate red block”, “pick red block”, “locate red bowl”, and “place red bowl”, whereas the plan in the second task is four times longer, as the agent has to move four blocks instead of one.
@@
TableTop Robotic Stand. To evaluate LERa in a real-world environment, we used the XArm6 robotic platform, which was equipped with XArm grippers and RealSense L515 and D435 sensors. It performed manipulation tasks with two primary actions: “pick” and “place”. The planner, based on an open-source LLM, generated a plan using object recognition data, while the fine-tuned VLM-based checker evaluated the success of each action execution. Replanning was triggered when the checker detected execution errors, such as misplacement or misclassification of objects. The experiments consisted of 18 different tasks involving the manipulation of various sets of objects, moving them from a table to a container. Notably, in this real-world setting, errors frequently occurred due to inaccuracies in the gripper position predicted by the agent. This led to failures in grasping or placing objects. Additionally, some episodes involved initial misclassification by the visual detector, resulting in an incorrect initial plan and, consequently, the inability to correctly execute the human instruction.